What is the exact value of this expression. sec^2 50°- tan^2 50°
Answer by nerdybill(7384)

One of the trigonometric identity is:
1+tan^2 u = sec^2 u
.
Here is a site w/trig identities:
http://www.sosmath.com/trig/Trig5/trig5/trig5.html
.
so, if we started with:
1+tan^2 u = sec^2 u
subtracting tan^2 u from both sides:
1 = sec^2 u - tan^2 u
.
Therefore:
sec^2 50°- tan^2 50° = 1